OpenStudy (anonymous):

What is the slope of a line that passes through 2,12 and -1,3 Hint: think rise over run

OpenStudy (anonymous):

2= x1 12=y1 -1=x2 3=y2 \[\frac{ y2-y1 }{ x2-x1 } \]

OpenStudy (anonymous):

(3-12)/(-1-2)=-9/-3=3 slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

\[\frac{ 3-12 }{ -1-2 }\]

OpenStudy (anonymous):

\[\frac{ -9 }{ -3 }\] = \[\frac{ 9 }{ 3 } = 3\]

OpenStudy (anonymous):

Therefore, the slope equals 3.